OpenStudy (anonymous):

What is the slope of the line that passes through the points (2, 8) and (6, 12)? A. –5 B. –1 C. 1 D. 5

OpenStudy (black_ninja123):

Formula \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

ok so @Black_ninja123 how can i use that in my problem?

OpenStudy (anonymous):

and sorreh if i seem dumb. im rlly bad at this topic

OpenStudy (black_ninja123):

y2 is 12 y1 is 6 x2 is 8 x1 is 2

OpenStudy (black_ninja123):

\[\frac{ 12-6 }{ 8-2 }\]

OpenStudy (black_ninja123):

\[\frac{ 6 }{ 6 } = 1\]

OpenStudy (anonymous):

ooh so its c??

OpenStudy (black_ninja123):

Yes

OpenStudy (black_ninja123):

Do you get it?

OpenStudy (anonymous):

omg thank you!!! you get medal for sure! and yes! i will use this for the others!! :D THANK U!

OpenStudy (black_ninja123):

Your welcome and thank you for the medal